Alcohol consumption slows people's reaction times. In a controlled government test, it takes a certain driver ?1=0.31 s to hit the brakes in a crisis when unimpaired and ?2=1.19 s when drunk. If the car is initially traveling at ?=83 km/h, how much farther does the car travel before the brakes are applied when the person is drunk than it travels when the person is sober?
Answer #1

Distance covered by a drunk person = speed * time

Distance covered by a drunk person = 23.055 m/s * 1.19 = 27.436 m

Now,

Distance covered by a sober person = 23.055 * 0.31 = 7.147 m/s

Therefore,

A drink person's car will travel about 20.3 meters more than a sober person's car